# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v1

@@ Appendix H Calculation of sequential measurement probabilities
$e^{iH\tau}=\sum_{m=0}^{\infty}e^{im\omega\tau}|m\rangle\langle m|\ ,$ (H5)
setting $\omega=1$ as always, we define the matrix elements
@@ I.1 Outline and Preliminaries
This appendix is devoted to proving that if Tsirelson’s inequality is violated by a given quantum state, with support entirely in $\mathcal{H}_{k}$, then that state will satisfy all four LG3s (Eqs 1.12-1.15) if the measurement times $t_{1},t_{2},t_{3}$ are the same as in the Tsirelson scheme. This result highlights how the Tsirelson test probes nonclassicality in a different way than the standard LG3s and begins to connect the two frameworks.
One simply needs to show two statements hold true. The conjunction of these two statements would imply that all four LG3s are greater than zero:
@@ I.3 $L_{1}$ is the smallest LG3 on $\mathcal{H}_{0},\mathcal{H}_{1}$ and $\mathcal{H}_{2}$
Recall from Appendix G that for $|\psi_{k}\rangle\in\mathcal{H}_{k}$, $C_{12}=C_{13}=C_{23}=C$. Therefore, on these subspaces, $L_{1}=1+3C$ and the rest are 12C1-2C. So, by showing that $C_{12}$ is negative we will have shown that $L_{1}$ is the smallest LG3. Note,
@@
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{\varepsilon}{2\pi}\ ,$ (K2)
@@
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{E_{L}}{2\pi}$ (K4)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}\ ,$ (K5)
@@
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{6}\left(\sum_{n=0}^{2}\left\langle Q\left(\frac{2\pi n}{N+1}\right)\right\rangle+3\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}$ (K6)

### v2

@@ Appendix H Calculation of sequential measurement probabilities
$e^{iH\tau}=\sum_{m=0}^{\infty}e^{im\tau}|m\rangle\langle m|\ ,$ (H5)
setting $\omega=\hbar=1$ as always, we define the matrix elements
@@ I.1 Outline and Preliminaries
This appendix is devoted to proving that if Tsirelson’s inequality is violated by a given quantum state, with support entirely in $\mathcal{H}_{k}$, then that state will satisfy all four LG3s (Eqs 1.12-1.15) if the measurement times $t_{1},t_{2},t_{3}$ are the same as in the Tsirelson scheme.
To show this claim, one simply needs to show two statements hold true. The conjunction of these two statements would imply that all four LG3s are greater than zero:
@@ I.3 $L_{1}$ is the smallest LG3 on $\mathcal{H}_{0},\mathcal{H}_{1}$ and $\mathcal{H}_{2}$
Recall from Appendix G that for $|\psi_{k}\rangle\in\mathcal{H}_{k}$, $C_{12}=C_{13}=C_{23}=C$. Therefore, on these subspaces, $L_{1}=1+3C$ and the rest are 1C1-C. So, by showing that $C_{12}$ is negative we will have shown that $L_{1}$ is the smallest LG3. Note,
@@
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{\varepsilon}{2\pi}\ ,$ (K2)
@@
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{E_{L}}{2\pi}$ (K4)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}\ ,$ (K5)
@@
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{6}\left(\sum_{n=0}^{2}\left\langle Q\left(\frac{2\pi n}{3}\right)\right\rangle+3\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}$ (K6)